### v1

# GraCoRe: Benchmarking Graph Comprehension and Complex Reasoning in Large Language Models

Abstract
Evaluating the graph comprehension and reasoning abilities of Large Language Models (LLMs) is challenging and often incomplete. Existing benchmarks focus primarily on pure graph understanding, lacking a comprehensive evaluation across all graph types and detailed capability definitions. This paper presents GraCoRe, a benchmark for systematically assessing LLMs’ graph comprehension and reasoning. GraCoRe uses a three-tier hierarchical taxonomy to categorize and test models on pure graph and heterogeneous graphs, subdividing capabilities into 10 distinct areas tested through 19 tasks. Our benchmark includes 11 datasets with 5,140 graphs of varying complexity. We evaluated three closed-source and seven open-source LLMs, conducting thorough analyses from both ability and task perspectives. Key findings reveal that semantic enrichment enhances reasoning performance, node ordering impacts task success, and the ability to process longer texts does not necessarily improve graph comprehension or reasoning. GraCoRe is open-sourced at https://github.com/ZIKEYUAN/GraCoRe.
GraCoRe: Benchmarking Graph Comprehension and Complex Reasoning in Large Language Models
Zike Yuan1,2, Ming Liu1, Hui Wang2,*, Bing Qin1,*, 1Harbin Institute of Technology, Shenzhen, China, 2Peng Cheng Laboratory, Shenzhen, China
footnotetext: * B. Qin and H. Wang are corresponding authors.
1 Introduction
Graph understanding and complex reasoning are critical capabilities of Large Language Models (LLMs) due to their wide-ranging applications in fields such as social network analysis, drug discovery, recommendation systems, and spatiotemporal predictionsBrown et al. (2020). Thus, the ability to understand and reason about graph-structured data is particularly essential for advancing Artificial General Intelligence (AGI)Zhao et al. (2023). Graph-structured data primarily includes homogeneous and heterogeneous graphs. Research on homogeneous graphs often focuses on graph-structure-specific issues, such as protein-protein interaction predictionRao et al. (2014).In contrast to homogeneous graphs, pure graphs are simpler and do not contain attribute information for nodes and edge, such as graph-theoretic problemsBorgatti and Everett (2006). For heterogeneous graphs, tasks often rely on their rich semantic information, such as those related to knowledge graph reasoning. However, the use of LLMs for parsing, understanding, and reasoning about graph-structured data remains a challenging area of study.
Figure 1: GraCoRe encompasses two overarching abilities and 19 distinct tasks within LLM on graph scenarios, facilitating a granular benchmarking from basic perceptivity to advanced interactivity.
Firstly, the reasoning and comprehension capabilities of current large language models significantly deteriorate when handling complex graph-structured data with numerous nodes and edges. This degradation is partly due to the models’ difficulty in processing long textual inputs, especially when graph-structured data is described through lengthy texts. Long texts not only increase the computational burden but also introduce noise and redundant information, further diminishing the models’ ability to capture critical details. Secondly, textual descriptions of graph-structured data often involve complex entity relationships and abstract concepts, requiring models to both understand explicit information and infer implicit connections and relationships. Current research tends to rely on direct mappings from graph structure to answersWang et al. (2024), overlooking the potential to enhance the models’ deeper reasoning capabilities. Given the lack of systematic definitions and evaluation standards for the capabilities of large language models in graph understanding and reasoning, it is imperative to design a benchmark that extensively tests these abilities based on the models’ capabilities.
In previous research, several benchmarks have been proposed to evaluate the understanding and reasoning capabilities of LLMs on graphs. However, existing benchmarks for evaluating the graph understanding capabilities of LLMs exhibit several limitations. Firstly, there is the issue of limited generalization: current benchmarks predominantly test either pure graphs or heterogeneous graphs in isolation, failing to provide a unified and systematic evaluation across both graph structures. Additionally, there is a lack of clear definition regarding model capabilities: traditional benchmarks for LLMs on graphs are primarily task-driven, inadequately assessing the specific abilities of LLMs on graph data. Hence, there is a need to develop better benchmarks that evaluate LLMs’ capabilities more comprehensively and from an ability-based perspective, specifically focusing on graph-structured data understanding and reasoning. Lastly, there is insufficient variety in model types and task categories: current benchmarks neither offer a clear classification of task types nor test a wide range of models.
To address these challenges, we propose the GraCoRe benchmark, as shown in Figure 1, which aims to explore and evaluate the understanding and reasoning capabilities of mainstream LLMs. We have designed a three-tier hierarchical ability taxonomy that includes both capability-based and dataset-based categories. This taxonomy meticulously defines the model’s capabilities and ensures greater generalization in the testing range. Regarding model types and task divisions, our benchmark tests multiple existing closed-source and open-source models, and divides tasks into multiple dimensions based on model capabilities.
Figure 2 illustrates the overall framework of our ability taxonomy. The first layer outlines two progressive overall capabilities: graph understanding and graph reasoning. Graph understanding is the most fundamental ability, reflecting the model’s proficiency in comprehending the nodes, edges, and overall structure of the graph within the contextual data. Built upon this foundation, graph reasoning encompasses the model’s ability to utilize graph-structured data to infer implicit information. The second layer categorizes the capabilities of LLMs into four types based on different data types. In the third layer, these four categories are further divided into 10 distinct capabilities, which are ultimately tested through 19 tasks. This taxonomy provides evaluation results across three levels, from general to detailed, allowing for the identification of deficiencies in models at varying levels of granularity. For each task, we meticulously design specific prompts and construct them based on predefined rules, ultimately textualizing the structural information. GraCoRe comprises a total of 11 datasets, consisting of 5,140 graphs. We control the complexity of these graphs by adjusting factors such as graph size and network sparsity.
Figure 2: Our three-tier hierarchical ability taxonomy.
Finally, we conducted extensive experiments on GraCoRe to evaluate the graph understanding and reasoning abilities of large language models. We tested three closed-source LLMs and seven open-source LLMs. Our findings include the following:
  • We identified graph reasoning as a critical weakness in current LLMs. Most models cannot balance both graph understanding and reasoning, with GPT-4o being the most comprehensive model in both aspects.
  • LLMs perform better on graph reasoning tasks enriched with semantic information compared to tasks involving purely structural graph reasoning, indicating that textual information can enhance the graph reasoning abilities of LLMs.
  • Models exhibit high sensitivity to the ordering of nodes in textual graph data. An ordered naming of nodes can significantly improve model performance on graph tasks.
  • The capability of models to handle longer text inputs does not affect their performance in graph understanding and reasoning tasks, regardless of whether the graphs are complex with long textual descriptions or simple with short descriptions.
2 Related Work
LLMs for Graph Recent advancements in LLMs for graph tasks includes several notable contributions. Li et al. (2023) categorizes these tasks into three types: Enhancer, Predictor, and Alignment. Pan et al. provides a forward-looking roadmap for the unification of LLMs and Knowledge Graphs (KGs).Chai et al. (2023) proposes an end-to-end method for solving graph-related problems. Furthermore, Das et al. (2023) investigates methods to improve the zero-shot reasoning ability of LLMs over structured data in a unified manner. Yao et al. (2024) explores the graph generation capabilities of LLMs through systematic task designs and extensive experiments.
Benchmarks for LLMs on Graph Most benchmarks for evaluating LLMs on graph tasks are based on task testing. NLGraphWang et al. (2024) introduced a simple test dataset for eight graph tasks, while GPT4GraphGuo et al. (2023) tested LLM capabilities on semantic tasks. Liu and Wu (2023) assessed the capabilities of four LLMs in graph data analysis. Fatemi et al. (2023) proposed a method for describing graph data in text form. Perozzi et al. (2024) designed a hint method specifically for graph tasks. GraphInstructLuo et al. (2024) provided diverse processes and steps for graph data generation. HiGPTTang et al. (2024) proposed an evaluation method for heterogeneous graphs, and VisionGraphLi et al. (2024) assessed the capabilities of LLMs on image graphs.
Task Abbr. Description
Node Number NN Calculate the total number of nodes in a graph.
Average Degree AD Calculate the average degree of the nodes in a graph.
Connectivity Test CT Determine if the graph is connected, meaning there is a path between any two nodes.
Matrix Similarity MS Evaluate the similarity between two adjacency matrices created from LLM and target graph.
Tree Recognition TR Identify if a graph is binary tree.
Bipartite Recognition BR Identify if a graph is bipartite, meaning its nodes can be divided into two disjoint sets such that no two nodes within the same set are adjacent.
Breadth First Search BFS Perform a breadth-first traversal starting from a given node.
Neighborhood Query NQ Query all nodes that are neighbors of a specified node.
Relationship Query RQ Query for specific relationships between nodes in the graph.
Relation Number RN Count the number of relationship types in the graph.
Subgraph Extraction SE Extract a subgraph based on specified criteria or nodes.
Shortest Path SP Find the shortest path between two nodes in a graph.
Maximum Flow MF Calculate the maximum flow in a flow network.
Eulerian Path EP Determine if there is a path that visits every edge exactly once.
Hamiltonian Cycle HC Determine if there is a cycle that visits every node exactly once and back start node.
Traveling Salesman Problem TSP Find the shortest possible route that visits each node exactly once and returns to the origin node.
Graph Coloring GC Assign colors to the nodes of the graph so that no two adjacent nodes share the same color.
Node Classification NC Classify nodes into predefined categories based on their attributes or graph structure.
Link Prediction LP Predict whether a link (edge) exists between two nodes in the graph.
Table 1: The 19 tasks for LLMs on graph within GraCoRe.
3 GraCoRe
This section begins by describing a three-tier hierarchical taxonomy for large language models (LLMs) on graph data. Next, we explain the methodology used to collect the dataset. Finally, we present an analysis of the dataset’s statistics.
3.1 Hierarchical Ability Taxonomy
After analyzing Bai et al. (2024) evaluation of the LLMs in the multi-round dialogue task, we have developed a hierarchical taxonomy for classifying LLMs capabilities on graphs, essential for their evaluation. The taxonomy is structured into three levels, encompassing 19 tasks within 10 sub-capabilities. Table 1 summarizes each task with a brief description. This section elaborates on the three levels and their corresponding tasks, with detailed examples provided in the Appendix A.
Benchmark Graph Type #Graph Average Node Average Edge #Node Type #Edge Type #Task
PureGra Bipartite Graph 460 18.52 48.9 1 1 5
Tree Structure Graph 460 19 18.49 1 1 5
Graph Traversal Graph 460 19 52.61 1 1 5
Shortest Path Graph 460 19 52.7 1 1 5
Max Flow Graph 460 19 105.72 1 1 1
Graph Coloring 460 19 52.7 1 1 1
Hamiltonian Graph 460 19 71.7 1 1 1
TSP Graph 460 19 193 1 1 1
Eulerian Graph 460 19.48 104.43 1 1 1
HeterGra IMDBText 500 27.21 30.32 3 4 4
ACMText 500 157.64 170.92 4 8 2
GraCoRe / 5140 / / / / 19
Table 2: Benchmark Graph Statistics
3.1.1 Graph understanding
Understanding graph structure necessitates large language models capable of accurately answering questions about the graph’s basic properties and reconstructing its structural information from extensive text descriptions. This involves two core capabilities:
Pure graph understanding: Pure graphs refer to graph data containing only structural information, representing the simplest form of graph dataJin et al. (2023). The structural information of such graphs can be encapsulated in an adjacency matrix. Consequently, research on pure graphs often emphasizes the structural information. Assessing the ability of large language models to understand pure graphs should thus focus on their capacity to comprehend structural information. To this end, I have identified four sub-capabilities:
  • Pure Graph Attribute Understanding: The most intuitive measure of understanding graph structure information is the correct comprehension of its basic attributes, such as the number of nodes, average degree, and performance on several sub-tasks related to node connectivity.
  • Pure Graph Memory: This requires the large language model to reconstruct the input graph structure data, testing its memory capacity. This is specifically evaluated by the similarity scoreLahitani et al. (2016) of the reconstructed matrix.
  • Pure Graph Recognition: For graph data with different structures, the model must be able to recognize and distinguish them. This study uses bipartite graphs and tree graphs to evaluate this capability.
  • Graph Traversal: Traversal is fundamental to solving many graph theory reasoning problems. The model’s performance in reasoning tasks is influenced by its traversal capability. This study primarily tests whether the model can traverse a graph using Breadth-First Search (BFS)Beamer et al. (2013).
Heterogeneous graph understanding: Unlike pure graphs, heterogeneous graphs often contain rich semantic information, with much of the data collected from real-world scenarios. Consequently, understanding heterogeneous graphs typically involves grasping their semantic information, whereas understanding their structural information is less critical. We have refined this into two sub-capabilities:
  • Graph QA and Querying: Given the rich semantic information in heterogeneous graphs, the ability of large language models to perform question-answering and querying is crucial. This capability includes three sub-tasks: querying neighbor nodes, answering questions about node relationships, and querying the number of relationships.
  • Subgraph Extraction: This pertains to the overall understanding of relationships and nodes within heterogeneous graphs, assessing the model’s ability to extract relevant subgraphs.
3.1.2 Graph reasoning
Based on the graph understanding capabilities of LLMs, their graph reasoning abilities are also worth exploring. This ability requires the model to infer hidden information from known graph data. It includes the following two core capabilities:
Graph structure reasoning: Graph structural information reasoning requires large language models to understand the nodes, edges, and their connections to infer the overall structural characteristics of the graph or the structural patterns of specific subgraphs. For example, the model should be able to identify cycles, paths, tree structures, and hierarchical structures within the graph, and use these structural features for further reasoning. Tasks related to structural reasoning are primarily focused on graph theory problems. We classify the complexity of these problems into two categories: Simple Graph Theory Problems Reasoning and Complex Graph Theory Problems Reasoning. The classification criterion is the time complexity of the corresponding algorithms. Simple graph theory problems are solvable in polynomial time, while complex graph theory problems are NP-complete, requiring significantly more time to solve. This classification tests the model’s capability in reasoning about graph theory problems. We have selected three representative problems for testing within each of these categories.
Graph semantic reasoning: Unlike structure-based reasoning, semantic information reasoning in graphs requires large language models to deeply understand the semantic meanings of nodes and edges, and to reason based on this semantic information. This involves modeling and reasoning about entities, relationships, and their interactions within the graph. Based on these tasks, we subdivide the semantic reasoning capabilities of large language models into Node Entity Reasoning and Link Relationship Reasoning. Corresponding tasks include node classification and link prediction. Previous studies have primarily addressed these two problems using graph neural networks (GNNs) such as GCNKipf and Welling (2016) and GraphSAGEHamilton et al. (2017). These methods typically require large structured graph data for training and cannot directly utilize graph data containing text information for inference. Consequently, investigating the use of large models for semantic reasoning is highly significant. This research will explore whether text enhancement impacts the performance of LLMs on these two tasks
3.2 Data Collection
We first divide the dataset into pure graphs and heterogeneous graphs to test the capabilities of large language models on these two types of datasets. For pure graphs, we customize unique data generation prompts based on the specific characteristics of each task, generating corresponding graph structure data using manually set rules. The scale of the graph is defined by the number of nodes and the sparsity of the network, ensuring that the generated data meets the specific needs of each task. For heterogeneous graph data, we use the ACMWang et al. (2019) and IMDBFu et al. (2020) datasets, converting them into text-based graph data. These are constructed according to a manually specified graph structure description framework, and unique prompts are designed for each task to build the dataset.
After generating the benchmark datasets, we also designed specific few-shot prompts for each task to test the model’s capabilities. These prompt datasets will be included in the benchmark data to provide additional testing options. Finally, we will provide a standard answer for each task and filter out graph data that does not meet the corresponding task requirements.
Model Pure Graph Heterogeneous Graph
Attribute
Understanding
Graph
Memory
Graph
Recognition
Grap
Traversaling
Graph QA
and Quering
Subgraph
Extraction
NN AD CT MS TR BR BFS NQ RQ RN SE
GPT-4o 75.012 94.339 51.752 73.886 98.639 52.789 73.136 80.298 68.642 77.347 83.649
GPT-4 74.898 82.494 70.146 73.82 75.85 32.023 72.455 89.205 68.39 80.095 84.717
GPT-3.5 75.012 62.358 46.51 70.897 23.7 79.558 79.869 57.03 34.508 57.926 59.261
Llama3-ins-8b 69.065 32.597 32.881 56.348 69.277 30.239 79.718 39.398 67.886 73.499 44.663
Llama2-7b-chat 22.521 50.661 66.429 48.442 41.668 30.239 50.593 38.489 2.263 44.552 37.542
Chatglm3-6b 30.527 45.627 69.002 47.246 38.6 30.239 36.447 16.131 27.832 19.024 20.275
Chatglm2-32k-7b 18.69 27.563 22.492 7.851 48.241 100 16.249 64.211 27.832 19.024 31.312
Vicuna-v1.5-16k 40.933 38.372 0 47.446 29.397 52.14 30.698 46.851 66.123 68.003 38.343
Qwen2-7b-ins 68.322 14.238 71.289 65.582 50.432 62.037 37.885 53.849 67.886 23.849 76.707
Vicuna-v1.5-7b 24.523 51.253 69.002 7.984 23.7 30.239 22.452 14.041 68.138 36.185 23.034
Table 3: Standardized performance of graph understanding
Model Graph Structure Reasoning Graph Semantic Reasoning Average $z$ Total Score
Simple Graph
Theory Problems
Complex Graph
Theory Problems
Node Entity
Reasoning
Link Relationship
Reasoning
SP MF EP HC TSP GC NC LP
GPT-4o 99.268 41.714 81.963 61.792 80.692 58.327 86.743 79.702 1.01 1419.69
GPT-4 68.526 26.463 82.776 65.548 63.373 49.208 85.286 73.2 0.793 1318.473
GPT-3.5 39.607 32.563 80.845 82.338 48.218 76.353 75.344 62.262 0.419 1144.161
Llama3-ins-8b 58.365 47.815 38.887 91.617 20.074 85.048 46.462 70.62 0.226 1054.461
Llama2-7b-chat 50.55 81.369 23.444 29.095 24.404 65.75 35.835 43.583 -0.347 787.429
Chatglm3-6b 11.992 81.369 55.853 28.653 59.043 64.689 41.663 28.62 -0.421 752.831
Chatglm2-32k-7b 19.287 81.369 35.331 28.653 48.218 23.335 25.293 26.556 -0.595 671.508
Vicuna-v1.5-16k 43.776 20.362 24.257 28.653 20.074 8.49 15.694 42.551 -0.616 662.165
Qwen2-7b-ins 52.113 60.016 51.18 54.06 91.517 37.544 47.233 67.318 0.223 1053.057
Vicuna-v1.5-7b 56.021 26.463 24.968 29.095 43.889 30.758 39.949 5.091 -0.691 626.783
Table 4: Standardized performance of graph reasoning
3.3 Data Statistics
Table 2 shows several key statistics of our GraCoRe benchmark. We categorized the dataset based on graph structure into two main datasets: PureGra and HeterGra. Each main dataset contains multiple sub-datasets used for corresponding task testing. In total, there are 19 tasks with up to 5,140 graphs. Detailed statistics for each task can be found in the Appendix A.
For pure graph data, the datasets include graphs with 8 to 30 nodes, with 20 test graphs per dataset. This design is intended to assess the impact of graph complexity on the performance of large language models. For heterogeneous graph data, we divided them into IMDBText and ACMText datasets. The ACMText dataset is more complex and extensive, containing more semantic information than the IMDBText dataset. Therefore, the ACMText dataset is primarily used for complex reasoning tasks, including node classification and edge prediction.
GraCoRe is the first benchmark specifically focused on the fine-grained understanding and reasoning capabilities of large language models on graph data.
3.4 Evaluation
This thesis evaluates the output of large language models (LLMs) using exact match accuracy. We focus on various output types, including boolean values (e.g., graph recognition tasks), integers (e.g., path lengths), floating-point numbers (e.g., similarity and average degree), and lists of nodes (e.g., paths). For tasks with multiple possible solutions (e.g., BFS), we verify whether the output constitutes a valid answer.
Since the metrics of different GraCoRe tasks are incomparable and differently sensitive, less experienced audiences cannot easily compare and interpret results, which is also prevalent in recent LLM benchmarks like KolaYu et al. (2023). Therefore, we utilized standardized global scoresDyck et al. (2005) to evaluate the performance of LLMs on each task.
Given a task set $T=\{t_{i}\}_{i=1}^{|T|}$ and an evaluated model set $M=\{m_{j}\}_{j=1}^{|M|}$, so $x_{ij}$ represents the performance of model $m_{j}$ on task $t_{i}$. Then the standardized score $z$ can be calculated as:
$z_{ij}=\frac{x_{ij}-\mu\left(x_{i1},\ldots,x_{i|M|}\right)}{\sigma\left(x_{i1},\ldots,x_{i|M|}\right)},$ (1)
where $\mu$(·) and $\sigma$(·) denote the mean and standard deviation.Next, we use the Min-Max scalingPatro and Sahu (2015) method to adjust the scores to the range of 0-100, making it easier to observe and compare the results. The final scores are presented as:
$s_{ij}=100\cdot\frac{z_{ij}-\min(z)}{\max(z)-\min(z)},$ (2)
where the functions max ($z$) and min ($z$) correspond to the maximum and minimum of all $z_{ij}$ scores.
Figure 3: Performance of various LLMs for second and third layer ability dimension.
4 Experiments
4.1 Experimental Setup
Based on the GraCoRe benchmark, we aim to investigate whether language models can understand and reason about graph structures through textual descriptions of graph data. Additionally, we explore whether designed prompts can enhance the models’ performance on graph-related tasks.
Models and Settings We evaluated a total of eight popular models on the GraCoRe benchmark, including three closed-source models and five open-source models. The closed-source models are: GPT-4o, GPT-4, and GPT-3.5Achiam et al. (2023). The open-source models are: LLama3-ins-8b, LLama2-7b-chatTouvron et al. (2023), Chatglm3-6b, Chatglm2-32k-7bDu et al. (2021), Vicuna-v1.5-7b, Vicuna-v1.5-16k-7bChiang et al. (2023) and Qwen2-7b-insBai et al. (2023). More details about these models can be found in the appendix.
Figure 4: Effect of Graph Size
4.2 Main Results
Task Dimensional Analysis Table 3 and Table 4 shows the performance of various large language models across 19 different tasks in our GraCoRe benchmark. Among all tasks, graph understanding tasks are generally less challenging for the models, while graph reasoning tasks are more difficult. This conclusion is reflected in the performance of each model. Additionally, closed-source models, such as those developed by OpenAI, consistently demonstrated superior performance. The GPT-4o model achieved the highest total score of 1419, ranking first. Among open-source models, Llama3-8b and Qwen2-7b-ins also performed exceptionally well, scoring 1054 and 1053 respectively, ranking fourth and fifth, just behind the OpenAI models. In contrast, Chatglm2-7b model performance was less satisfactory compared to other open-source models. Regarding task types, performance on reasoning tasks was not ideal, but this outcome was expected.
Additionally, the average $z$-scores indicate a significant gap between open-source models and commercial closed-source models. Only the Llama3-8b and Qwen2-7b-ins models have $z$-scores greater than 0, indicating performance above the average level. This suggests that the open-source community needs more collaboration to support and improve the latest open-source large language models. It is noteworthy that these two open-source models have graph understanding and reasoning capabilities approaching those of the GPT-3.5 model.
Ability Dimensional Analysis We further analyze Table 3 and Table 4 to evaluate the overall performance of the models from a capability perspective. We use radar charts to present the second and third layers of the three-tier taxonomy in a more intuitive manner. Left part of Figure 3 shows the performance of LLMs across four different capability dimensions at the second layer, where each dimension’s score is the average of its respective tasks. It can be observed that most models perform well in graph understanding and semantic reasoning but need improvement in structural reasoning tasks. For most open-source models, performance in heterogeneous graph understanding tasks is not very satisfactory.
Right part of Figure 3 provides a more granular view of LLMs’ performance across ten capability dimensions at the third layer. It is evident that the reasoning abilities of large language models in graph theory problems are not very strong, particularly in complex graph theory tasks. Furthermore, GPT-4 and GPT-4o exhibit robust and balanced graph processing capabilities, especially excelling in complex tasks. In contrast, other models show less balanced performance in graph processing. For instance, the Vicuna-v1.5-16k-7b model excels in understanding heterogeneous graph structures but performs poorly in other areas, suggesting that the rich semantic information may enhance the model’s graph processing capabilities.
Long-Text-Specific Models Since graph structure data described in text often consists of long texts, the ability of models to handle long texts is also worth noting. As shown in Table 3, models capable of processing long texts, such as Chatglm2-32k-7b and Vicuna-v1.5-16k-7b, performed poorly in graph processing tasks. Compared to other models, their performance was even lower. This suggests that despite being designed for long text input and output, these models still require further development and training to effectively enhance their graph processing capabilities.
4.3 Further Analysis
Effect of Graph Size Figure 4 illustrates the performance of the model across four tasks as the number of nodes increases. Specifically, we used two graph understanding tasks and two graph reasoning tasks to evaluate the model’s performance. The results indicate that as the number of nodes increases, the performance on each task decreases. In the graph understanding tasks, the impact is particularly significant on the average degree task. We hypothesize that the increase in nodes leads to a substantial rise in computational demands, resulting in decreased accuracy. In the graph reasoning tasks, we observed that for the shortest path problem, the performance of GPT-3.5 diminishes with an increasing number of nodes. However, for the other three open-source models, the node size appears to have minimal impact. This may be because the added nodes do not affect the search for the original path, though further experiments are needed to confirm this hypothesis.
Effect of Random Sort Since our data consists of randomly generated graphs with nodes named by numbers, the performance of node-related tasks may be affected by changes in node order. In this study, we examine the impact of text order on the understanding of graph structures by large language models by comparing random sorting with sequential sorting. The results in Table 5 indicate that the model’s performance under sequential sorting is often superior to that under random sorting, particularly in graph path reasoning tasks, where the impact is significant. This suggests that renaming nodes and ordering them sequentially can enhance model performance in path reasoning problems. However, it also highlights the model’s lack of training on graph data with random sorting.
Model NN ST BR SP HP GPT-3.5 sort 0.993 0.954 0.504 0.117 0.243 random 0.968 0.865 0.346 0.070 0.048 Qwen2-7b sort 0.876 0.874 0.396 0.165 0.115 random 0.773 0.885 0.374 0.159 0.000 Llama3-ins-6b sort 0.889 0.735 0.200 0.189 0.285 random 0.786 0.809 0.200 0.198 0.059
Table 5: Performance of different models with sorted and random sorted grpah text input.
Effect of Text Enhancement
Figure 5: Effect of Text Enhancement
Heterogeneous graphs often contain rich semantic information, which typically aids models in understanding and reasoning about graph text information. If enhanced text information, such as paper titles and abstracts in the ACMText data is removed, can large language models perform reasoning based solely on the remaining structural information? Figure 5 compares the performance of different large language models on the node classification problem. The results indicate that large language models can reason about graphs without enhanced text information, particularly the GPT-4 and GPT-4o models, which maintain strong node prediction capabilities even without text information. However, for the GPT-3.5 model, the absence of text information significantly impacts performance. For other open-source models, the presence or absence of text information appears to have minimal impact.
5 Conclusion
This paper introduces GraCoRe, a benchmark designed to evaluate the capability of large language models (LLMs) in understanding and reasoning with graph-structured data. We develop a detailed, multi-level classification system to assess model performance on graph-based tasks. Utilizing GraCoRe, we evaluate 10 prominent LLMs, revealing significant deficiencies in their graph reasoning abilities. Extensive experimental results demonstrate the effectiveness of our benchmark in measuring the performance of LLMs on graph tasks.
6 Limitations
As LLMs continue to develop, the volume of training data and their capacity to represent graphs are likely to increase. Our current evaluation may not encompass all their capabilities, and some models might incorporate our data for training, potentially influencing the final evaluation outcomes. In the future, we aim to continually refine and update the GraCoRe benchmark to more effectively assess the graph understanding and reasoning abilities of emerging LLMs.
Appendix A Details on the Data Generation
Our dataset mainly includes two types of graph data, heterogeneous graph data and pure graph data. In addition, based on these two types of data, we set up 19 tasks to evaluate the large model, which includes a total of 11 datasets. This diversity ensures the comprehensiveness and stability of our data.
Figures 6 to 26 illustrate examples of prompts and graph data descriptions for each task. To ensure that the generated results align with our capabilities and task requirements, we concatenate the prompts for each task with the initial prompts
Figure 6: The initial instructions for answer generation.
Figure 7: The initial instructions for answer generation.
Figure 8: The unique prompt for the Node Number task.
Figure 9: The unique prompt for the Average Degree task.
Figure 10: The unique prompt for the Connectivity Test task.
Figure 11: The unique prompt for the Matrix Similarity task.
Figure 12: The unique prompt for the Shortest Path task.
Figure 13: The unique prompt for the Bipartite Recognition task.
Figure 14: The unique prompt for the Eulerian Path task.
Figure 15: The unique prompt for the Graph Coloring task.
Figure 16: The unique prompt for the Breadth First Search task.
Figure 17: The unique prompt for the Hamiltonian Cycle task.
Figure 18: The unique prompt for the Maximum Flow task.
Figure 19: The unique prompt for the Tree Recognition task.
Figure 20: The unique prompt for the TSP task.
Figure 21: The unique prompt for the Relation Number task.
Figure 22: The unique prompt for the Neighborhood Query task.
Figure 23: The unique prompt for the Relationship Query task.
Figure 24: The unique prompt for the Subgraph Extraction task.
Figure 25: The unique prompt for the Node Classification task.
Figure 26: The unique prompt for the Link Prediction task.